# FloorplanQA: A Benchmark for Spatial Reasoning in LLMs using Structured Representations

### v2

@@ E.2 Token-Limit Analysis and Valid-Only Accuracy
Question Room [Uncaptioned image] gpt-5 [Uncaptioned image] gpt-oss 120b [Uncaptioned image] DeepSeek R1-0528 [Uncaptioned image] Gemini Flash 2.5 [Uncaptioned image] gpt-5 mini-2025 [Uncaptioned image] gpt-oss 20b [Uncaptioned image] Qwen3 30B Think.
@@
Appendix F Additional Ablations and Analysis
To further verify the robustness and generalizability of our evaluation setup, we conducted supplementary analyses beyond the main accuracy tables.
@@ Appendix F Additional Ablations and Analysis
Figure 5: Radar plots for general models, showing mean and standard deviation of accuracy across (top) models and (bottom) question type.
@@
Figure 6: Radar plots for reasoning models, showing mean and standard deviation of accuracy across (top) models and (bottom) question type.
Prompt Sensitivity Ablation
@@
To assess model robustness to prompt variation, we conducted an ablation in which each question was regenerated using the same template but with alternate object references. For example, a question originally referring to a ”sofa” might instead use a ”bookshelf” in the regenerated version. This allowed us to evaluate whether model performance remains stable under changes in object content while preserving linguistic structure.
@@
As shown in Table 9, accuracy under prompt regeneration is broadly stable; larger models change little, and smaller ones fluctuate modestly. We observe somewhat higher sensitivity for Repositioning: paraphrases can implicitly select different target objects or motion directions, occasionally introducing additional complexity or non-movable cases. Overall, the evaluation appears robust to prompt-level variations.
@@
Appendix G Case Studies by Question Type
@@ G.3 Free Space
We compute the free floor area using geometric operations provided by Shapely (Gillies et al., 2007). All object polygons within the room are merged using unary_union to correctly handle overlaps. The occupied area is then obtained from this union, and the free area is computed as
@@ G.4 View Angle
(3) Let the global north vector be $\mathbf{n}=(0,1)$ (already unit length). Compute the cosine via $\cos\theta=\mathrm{clip}\!\big(\hat{\mathbf{d}}\cdot\mathbf{n},\,-1,\,1\big)$.
(4) Convert to degrees and take the smallest absolute angle: $\theta=\arccos(\cos\theta)\cdot\dfrac{180}{\pi}\ \in[0^{\circ},180^{\circ}]$.
@@ G.8 Max Box
Let $R$ be the room polygon and $O$ the union of all object polygons except rugs (soft coverings). Compute free space $F=R\setminus O$. Search over orientations $\theta\in[0,\pi)$: rotate $F$ by $\theta$, find the largest axis-aligned empty rectangle inside the rotated $F$, record $(w_{\theta},h_{\theta})$ and area $A_{\theta}=w_{\theta}h_{\theta}$, then map back to get $(w^{*},h^{*},\theta^{*})$ with $A^{*}=\max_{\theta}A_{\theta}$.
@@ Appendix H Prompts
Appendix I Supplementary Accuracy Analyses
@@
Question Room [Uncaptioned image] claude sonnet-4 [Uncaptioned image] gpt-4.1 [Uncaptioned image] Kimi-K2 Instruct [Uncaptioned image] Qwen3 Coder-480B [Uncaptioned image] Qwen3 235B [Uncaptioned image] gpt-4.1 mini [Uncaptioned image] Qwen3 30B [Uncaptioned image] Devstral Small
@@
Question Room [Uncaptioned image] claude sonnet-4 [Uncaptioned image] gpt-4.1 [Uncaptioned image] Kimi-K2 Instruct [Uncaptioned image] Qwen3 Coder-480B [Uncaptioned image] Qwen3 235B [Uncaptioned image] gpt-4.1 mini [Uncaptioned image] Qwen3 30B [Uncaptioned image] Devstral Small
@@
Question Room [Uncaptioned image] gpt-5 [Uncaptioned image] gpt-oss 120b [Uncaptioned image] DeepSeek R1-0528 [Uncaptioned image] Gemini Flash 2.5 [Uncaptioned image] gpt-5 mini-2025 [Uncaptioned image] gpt-oss 20b [Uncaptioned image] Qwen3 30B Think.
@@
Question Room [Uncaptioned image] gpt-5 [Uncaptioned image] gpt-oss 120b [Uncaptioned image] DeepSeek R1-0528 [Uncaptioned image] Gemini Flash 2.5 [Uncaptioned image] gpt-5 mini-2025 [Uncaptioned image] gpt-oss 20b [Uncaptioned image] Qwen3 30B Think.
Pair distance K 99.8 99.3 98.0 96.3 100.0 94.2 97.7
LR 98.8 99.3 99.0 96.0 99.7 93.5 97.5
B 98.3 99.5 96.8 95.5 99.7 93.8 95.3
HSSD 69.0 78.5 25.5 12.5 32.5 40.5 18.0
Placement K 84.7 92.0 89.0 59.7 90.8 85.7 68.2
LR 75.5 89.0 82.2 53.3 86.3 78.5 46.5
B 61.2 83.5 72.0 35.8 81.2 62.0 34.5
HSSD 70.0 85.0 79.0 16.0 75.5 74.5 28.5
Reposi- tioning K 83.0 85.5 79.2 90.5 84.5 70.8 61.5
LR 85.5 89.8 86.2 91.2 92.8 87.8 77.0
B 77.8 83.3 83.3 85.2 84.3 78.0 69.2
HSSD 49.5 60.5 47.5 18.5 53.5 40.0 27.0
Free space K 82.5 99.0 93.0 93.3 99.5 94.8 97.0
LR 47.0 83.3 18.3 17.7 78.5 53.2 0.0
B 50.5 87.5 34.8 33.3 82.2 74.0 1.0
HSSD 19.5 31.0 6.5 1.0 5.0 9.0 1.0
Visibility K 94.8 94.2 71.3 26.8 98.0 91.5 78.8
LR 95.2 94.0 52.0 11.3 98.0 89.3 70.8
B 94.2 92.5 53.5 11.2 95.5 89.2 64.2
HSSD 57.0 70.0 10.0 0.5 39.0 45.5 3.0
View Angle K 96.2 98.5 73.7 92.5 88.3 93.5 98.5
LR 93.3 97.3 68.2 91.5 84.5 92.0 98.3
B 95.2 98.2 75.2 93.8 86.8 91.3 98.3
HSSD 59.5 74.0 13.5 20.0 25.5 37.5 26.0
Max Box K 48.5 62.8 50.5 3.7 85.2 31.3 16.7
LR 17.3 28.2 8.0 0.0 60.3 3.8 0.8
B 13.0 30.3 11.0 0.0 61.2 5.8 0.5
HSSD 5.0 9.5 2.5 0.0 17.0 0.5 0.0
Shortest path (valid) K 64.7 64.2 12.3 3.2 52.3 43.0 14.2
LR 21.2 66.0 12.5 1.5 53.7 37.7 16.7
B 58.2 57.8 7.8 1.0 52.0 30.0 14.3
HSSD 28.5 33.5 8.5 0.0 16.5 13.5 1.0
Shortest path (Fréchet) K 56.3 55.5 11.5 3.2 50.5 42.2 14.0
LR 12.3 40.5 9.3 1.3 47.5 28.5 16.3
B 39.3 44.8 6.0 0.8 47.7 24.2 14.3
HSSD 22.5 26.5 2.5 0.0 12.0 14.0 1.0

### v3

@@ E.2 Token-Limit Analysis and Valid-Only Accuracy
Question Room [Uncaptioned image] GPT-5
[Uncaptioned image] gpt-oss
120b
[Uncaptioned image] DeepSeek
R1-0528
[Uncaptioned image] Gemini
Flash 2.5
[Uncaptioned image] GPT-5
mini-2025
[Uncaptioned image] gpt-oss
20b
[Uncaptioned image] Qwen3
30B Think.
@@
E.3 Aggregated Truncation Rates
Table 7 reports aggregated truncation and error-type statistics for each model, computed by averaging outcomes across all room types and all question categories in the full dataset. A response is counted as truncated when it terminates with the TOKEN LIMIT stop reason. We additionally report the fraction of invalid-format responses that could not be parsed into a valid answer under our evaluation protocol.
For completeness, Table 7 also includes the proportions of wrong and correct responses, as well as an alternative accuracy (% alt), computed by extracting the last numeric value or the final list from each model’s output. This alternative metric serves purely as a robustness check on the parsing and evaluation pipeline.
Table 7: Error-type distribution and accuracy on the full dataset. Values are aggregated for each model by averaging across all room types and all questions. Truncation (% trunc.) quantifies the fraction of responses that were terminated due to the TOKEN LIMIT stop reason. Invalid-format (% invalid) denotes responses that could not be parsed into a valid answer. Alternative accuracy (% alt) reports accuracy computed using the last numeric value or the last list in the model’s output.
Model % trunc. % invalid % wrong % correct % alt
gpt-5 3.86 0.27 30.42 65.45 65.45
gpt-oss-120b 2.05 1.05 21.41 75.49 75.53
DeepSeek-R1 31.31 0.78 17.26 50.65 50.65
gpt-5-mini 15.50 1.34 12.54 70.62 70.62
Gemini Flash 2.5 50.58 0.13 10.14 39.15 39.15
gpt-oss-20b 16.94 1.18 21.70 60.18 60.31
Qwen3-30B Think 38.95 0.09 16.52 44.44 44.44
claude-sonnet-4 0.05 0.07 41.09 58.80 58.80
gpt-4.1 0.14 0.09 41.52 58.25 58.25
Kimi-K2 0.15 0.03 52.04 47.78 47.78
Qwen Coder-480B 0.02 0.09 49.40 50.48 50.48
Qwen-235B 13.64 0.13 20.31 65.92 65.98
gpt-4.1-mini 0.67 0.10 42.18 57.05 57.05
Qwen-30B Instr 6.85 3.09 38.47 51.59 51.84
Devstral Small 2.75 1.56 65.64 30.05 30.07
E.4 Radar summaries
For readability, we also provide radar visualizations that summarize accuracy and variance across models and question types, complementing the main tables.
@@ E.4 Radar summaries
Figure 6: Radar plots for general models. Top: mean accuracy (left) and standard deviation (right) across models, grouped by room type. Bottom: mean accuracy (left) and standard deviation (right) across question types.
@@
Figure 7: Radar plots for reasoning models. Top: mean accuracy (left) and standard deviation (right) across models, grouped by room type. Bottom: mean accuracy (left) and standard deviation (right) across question types.
Appendix F Tools and VLM experiments
To complement the text-only benchmark, we additionally evaluate (i) tool-augmented reasoning using an integrated Python interpreter and (ii) vision-language models (VLMs) using rendered floorplan images. These studies measure the potential benefit of external computation and visual input while keeping the task definitions and scoring identical to the main benchmark.
Tool-augmented setting.
We enable the Code Interpreter tool for GPT-4.1 and GPT-4.1-mini, allowing the model to invoke a Python sandbox during inference. The model is instructed to use Python whenever numeric computation is needed (e.g., distances, angles, polygon centroids or areas), and to output a final numeric answer in the same format as the raw setting. While the interpreter improves numeric precision, models can still fail on complex tasks either due to incorrect spatial reasoning or because the generated code is incomplete or erroneous. Results are reported in Table 8.
VLM setting and renderings.
We also evaluate VLM inputs by providing a top-down rendered floorplan image (720×720 pixels) alongside both the text question and the complete JSON layout. This multimodal setting allows models to leverage visual rendering while retaining access to precise symbolic coordinates. We use two controlled rendering styles derived from the same symbolic layout: (i) a minimal contour-based rendering, where room and object polygons are shown as boxes with text labels (similar to Fig. 4, right), and (ii) an icon-based rendering, where objects are depicted with furniture icons instead of bare contours (similar to Fig. 2). Because HSSD contains a wider variety of object categories than our current icon library, icon-based renderings are used only for synthetic layouts, while HSSD layouts use the contour style. VLM results under these input conditions are summarized in Table 8.
Table 8 shows that Code Interpreter substantially improves metric-dominated scalar tasks (distance, angle, visibility, placement), but offers limited benefit on planning-heavy tasks (Max Box, Shortest Path), indicating that remaining errors are primarily spatial rather than arithmetic. VLM inputs provide selective gains, mainly on generated layouts and for object-fit or metric cues, while not consistently improving global performance.
Table 8: Task accuracy for gpt-4.1-2025-04-14 and gpt-4.1-mini-2025-04-14 under three settings: raw text-only input (Raw), tool-augmented reasoning with a Python interpreter (Tools), and vision-language input using rendered floorplans (Img; Icons). HSSD columns report performance on 200 human-designed scenes. Generated columns report performance on 200 synthetic scenes (50 kitchens, 75 living rooms, and 75 bedrooms). Cell colors indicate gains (green) or drops (red) relative to the corresponding Raw setting for the same model and task.
HSSD Generated
Question GPT-4.1 GPT-4.1-mini GPT-4.1 GPT-4.1-mini
Raw Tools Img Raw Tools Img Raw Tools Img Icons Raw Tools Img Icons
Pair distance 56 99 60.5 37 98 51.5 95.5 99.5 99 99.5 86.5 98.5 94.5 99.5
Placement 64.5 95 70.5 76.5 87.5 78.5 65 92.5 61.5 71.5 76 95.5 70.5 81
Repositioning 47 83.5 44.5 40.5 68 37 56.5 48 42 39 70 59.5 45 50.5
Free space 16 44 21.5 15.5 42 26 41 28.5 55 45 24.5 27.5 26 22.5
Visibility 46.5 86.5 36.5 52 70 37 80 89 61.5 56.5 88 71.5 72.5 68
View angle 55 96 63.5 46 93 54 92 99 93.5 95.5 91.5 97 92.5 92.5
Max box 7.5 3 4 4.5 3.5 4 12 11.5 12 10.5 11 8 7.5 9.5
Shortest path (valid) 25 12.5 22.5 15 14 16.5 51 34 46 52 46.5 24.5 44 45
Shortest path (Fréchet) 22.5 12.5 23.5 12.5 9.5 13.5 40.5 31 44 45 24.5 21 33 27
As illustrative failure cases, we also observe that tool augmentation does not guarantee correctness for tasks that require precise geometric reasoning. In Max Box (Fig. 8(a)), the model-generated Python program performs an approximate rotated/grid search and returns a suboptimal rectangle (red, $5.57\,\mathrm{m}^{2}$), despite a larger valid axis-aligned solution existing (green, $7.57\,\mathrm{m}^{2}$). In a second example from Repositioning ( Fig. 8(b)), the ground-truth maximum downward displacement of the dishwasher is $1.96$ m, but the model’s tool-based code predicts zero movement. Inspection shows that the program treats a shared boundary (the dishwasher touching another object or wall) as a collision, and therefore incorrectly concludes that no valid motion is possible for this concrete task. These cases highlight that tools reduce arithmetic imprecision, but complex tasks may still fail due to imperfect model-written algorithms and sensitivity to geometric edge cases (e.g., boundary contact vs. overlap).
(a) Max Box tool failure. The model-written code returns a suboptimal rotated rectangle (red, $5.57\,\mathrm{m}^{2}$) despite a larger valid axis-aligned rectangle existing (green, $7.57\,\mathrm{m}^{2}$).
(b) Repositioning tool failure. The true maximum downward move is $1.96$ m, but the model’s code outputs 0 by treating boundary contact as collision.
Figure 8: Representative failure cases in the tool-augmented setting. While the python interpreter improves numeric precision, complex tasks can still fail due to imperfect model-written code and geometric edge cases.
Appendix G Sensitivity studies for numeric tolerances and path thresholds
G.1 Sensitivity studies for numeric tolerances.
We examine the robustness of our numeric evaluation tolerances for both scalar and area-based question types. For clarity and cost control, we conduct these sensitivity analyses on a representative subset of six models (three reasoning-focused and three general-purpose).
For scalar metrics (e.g., Pair distance, View angle, Repositioning, and Max Box), Figure 9 shows that relative errors are sharply concentrated near zero across models, with a clear knee before $\varepsilon_{\mathrm{rel}}=2\%$ (red dashed line). The accompanying tolerance sweep in Figure 11 (left) demonstrates smooth, monotone accuracy gains while changing tolerance from $0.5\%$ to $5\%$ without changing model rankings, indicating that the selection of $2\%$ lies in a stable regime rather than being outcome-sensitive.
Figure 9: Aggregated relative-error distributions for scalar metrics across models. Left: CDF of relative error; the dashed line marks $\varepsilon_{\mathrm{rel}}=2\%$. Right: Histogram showing a strong peak near zero and a thin long tail.
For free-space (area) questions, Figure 10 shows a broader, heavier-tailed error distribution; nevertheless, $5\%$ tolerance lies near saturation for higher-performing models while remaining strict for weaker ones. The sweep in Figure 11 (right) confirms that relaxing area tolerance from $1\%$ to $10\%$ yields gradual improvements and preserves qualitative conclusions.
Figure 10: Aggregated relative-error distributions for free-space (area) questions across models. Left: CDF of relative error; the dashed line marks $\varepsilon_{\mathrm{rel}}=5\%$. Right: Histogram showing broader, heavier-tailed errors compared to scalar tasks.
Overall, these results justify our use of a $2\%$ tolerance for scalar outputs and a $5\%$ tolerance for complex area computations.
Figure 11: Tolerance sweeps for scalar (left) and free-space area (right) tasks. Aggregated accuracy increases smoothly as tolerance is relaxed (0.5–5% for scalars; 1–10% for areas), while model rankings remain unchanged.
G.2 Sensitivity studies for path thresholds
We analyze robustness of shortest-path scoring with respect to the Fréchet threshold $\tau$. Figure 12 shows that accuracy increases smoothly as $\tau$ is relaxed, and model rankings remain almost stable across the sweep. We therefore choose $\tau=0.6$ m as deviations within roughly 0.6 m correspond to paths that remain traversable for a person and allow minor alternate routes without accepting qualitatively different solutions.
In addition, path validity requires collision-free traversal under a clearance buffer of 0.15 m. This value represents a minimal safety margin; larger buffers (e.g., 0.3 m) would incorrectly invalidate many feasible paths in compact rooms, particularly around $\sim$0.6 m-scale kitchen utilities, and would over-penalize narrow but realistic layouts.
Figure 12: Shortest-path accuracy as a function of Fréchet tolerance $\tau$. Aggregated accuracy rises smoothly with increasing tolerance, while model ordering changes slightly. We use $\tau=0.6$ m as a moderate, human-scale tolerance.
Appendix H Ablations
H.1 Format Sensitivity Ablation
To evaluate the robustness of layout interpretation under alternate encodings, we replaced the standard JSON layout with a semantically equivalent XML version. This preserves all geometric and object-level content while changing only syntax. We select three tasks that show variance across HSSD layouts: View Angle, Visibility, and Repositioning. These span different reasoning categories (Metric, Topology, Dynamic) and have moderate baseline accuracy, where format changes are more meaningful than on near-floor tasks like Max Box. We test three models: two reasoning (gpt-oss-120B, gpt-oss-20B) and one general (Qwen3-235B-A22B). As shown in Table 9, accuracy stays stable across formats ($\pm$3 points), suggesting models encode layout semantics independently of serialization syntax.
Model Repositioning View Angle Visibility
gpt-oss-120B 60.5 → 59.0 74.0 → 74.0 70.0 → 72.0
gpt-oss-20B 40.0 → 39.0 37.5 → 38.5 45.5 → 43.5
Qwen3-235B-A22B 39.0 → 42.0 50.5 → 54.0 70.5 → 65.5
H.2 Object Reference Substitution Ablation
@@
To assess model robustness to object reference substitution, we conducted an ablation in which each question was regenerated using the same template but with different object references. For example, a question originally referring to a ”sofa” might instead refer to a ”bookshelf” in the regenerated version. This allowed us to evaluate whether model performance remains stable when object content changes while linguistic structure is preserved.
@@
As shown in Table 10, accuracy under object reference substitution is broadly stable; larger models change little, and smaller ones fluctuate modestly. We observe somewhat higher sensitivity for Repositioning: substitutions can implicitly select different target objects or motion directions, occasionally introducing additional complexity or non-movable cases. Overall, the evaluation appears robust to object reference substitution.
@@ H.3 Semantic Ablation
Table 11: Accuracy under semantic ablation. Each cell displays performance on the original prompt, followed by a version with permuted object labels (same geometry, different semantics).
@@
Appendix I Case Studies by Question Type
@@ I.3 Free Space
We compute the free floor area using geometric operations provided by Shapely (Gillies and others, 2007). All object polygons within the room are merged using unary_union to correctly handle overlaps. The occupied area is then obtained from this union, and the free area is computed as
@@ I.4 View Angle
(3) Let the global north vector be $\mathbf{n}=(0,1)$ (already unit length). Compute the clipped dot product: $c=\mathrm{clip}\!\big(\hat{\mathbf{d}}\cdot\mathbf{n},\,-1,\,1\big)$.
(4) Convert to degrees: $\theta=\arccos(c)\cdot\dfrac{180}{\pi}\ \in[0^{\circ},180^{\circ}]$.
@@ I.8 Max Box
Let $R$ be the room polygon and $O$ the union of all object polygons except rugs (soft coverings). Compute free space $F=R\setminus O$. Search over orientations $\theta\in[0,\pi)$: rotate $F$ by $\theta$, find the largest axis-aligned empty rectangle inside the rotated $F$, record $(w_{\theta},h_{\theta})$ and area $A_{\theta}=w_{\theta}h_{\theta}$, then map back to get $(w^{*},h^{*},\theta^{*})$ with $A_{max}=\max_{\theta}A_{\theta}$.
@@
Appendix K Supplementary Accuracy Analyses
@@
Question Room [Uncaptioned image] Claude
Sonnet-4
[Uncaptioned image] GPT-4.1
[Uncaptioned image] Kimi-K2
Instruct
[Uncaptioned image] Qwen3
Coder-480B
[Uncaptioned image] Qwen3
235B
[Uncaptioned image] GPT-4.1
mini
[Uncaptioned image] Qwen3
30B
[Uncaptioned image] Devstral
Small
@@
Question Room [Uncaptioned image] Claude
Sonnet-4
[Uncaptioned image] GPT-4.1
[Uncaptioned image] Kimi-K2
Instruct
[Uncaptioned image] Qwen3
Coder-480B
[Uncaptioned image] Qwen3
235B
[Uncaptioned image] GPT-4.1
mini
[Uncaptioned image] Qwen3
30B
[Uncaptioned image] Devstral
Small
@@
Question Room [Uncaptioned image] GPT-5
[Uncaptioned image] gpt-oss
120b
[Uncaptioned image] DeepSeek
R1-0528
[Uncaptioned image] Gemini
Flash 2.5
[Uncaptioned image] GPT-5
mini-2025
[Uncaptioned image] gpt-oss
20b
[Uncaptioned image] Qwen3
30B Think.
@@
Question Room [Uncaptioned image] GPT-5
[Uncaptioned image] gpt-oss
120b
[Uncaptioned image] DeepSeek
R1-0528
[Uncaptioned image] Gemini
Flash 2.5
[Uncaptioned image] GPT-5
mini-2025
[Uncaptioned image] gpt-oss
20b
[Uncaptioned image] Qwen3
30B Think.
Pair distance K 99.8 98.0 99.5 99.7 100.0 94.8 99.7
LR 98.8 99.3 99.8 100.0 99.8 94.1 99.8
B 98.3 99.5 99.3 99.8 99.8 94.5 100.0
HSSD 69.0 88.2 85.0 92.6 98.5 66.9 64.3
Placement K 97.5 92.0 93.8 98.1 97.2 89.7 96.7
LR 97.6 89.2 88.5 98.5 96.3 89.5 97.9
B 95.8 83.6 78.4 97.7 93.1 83.4 97.2
HSSD 95.9 85.0 84.0 100.0 90.4 87.7 96.6
Reposi- tioning K 83.3 85.5 86.4 91.4 84.5 71.4 66.2
LR 85.5 89.8 93.2 93.2 92.8 88.3 82.5
B 78.0 83.3 86.1 86.6 84.3 78.7 75.5
HSSD 50.5 60.8 71.4 77.1 74.8 51.6 73.0
Free space K 83.5 99.0 98.4 95.2 99.5 95.6 99.3
LR 47.1 83.5 33.0 30.3 79.8 61.6 0.0
B 50.6 87.5 39.7 48.1 83.0 80.3 1.2
HSSD 20.5 43.4 31.7 28.6 33.3 40.0 66.7
Visibility K 94.8 95.6 97.3 98.8 98.0 91.8 97.5
LR 95.2 95.6 97.8 95.8 98.3 90.2 96.2
B 94.2 93.6 96.4 95.7 97.0 89.3 96.2
HSSD 57.0 73.7 76.9 100.0 89.7 64.1 75.0
View angle K 96.2 98.5 98.9 97.9 99.4 94.4 99.5
LR 93.3 97.3 98.5 96.3 98.6 93.2 99.0
B 95.2 98.2 98.7 97.4 99.4 92.4 99.2
HSSD 59.5 80.0 87.1 81.6 94.4 60.5 76.5
Max box K 48.5 62.8 72.3 84.6 86.9 44.0 44.6
LR 17.3 28.2 21.9 0.0 64.0 9.8 41.7
B 13.0 30.3 27.1 0.0 64.2 13.3 27.3
HSSD 5.0 9.5 6.0 0.0 21.9 0.8 0.0
Shortest path (valid) K 64.8 64.5 57.4 82.6 68.4 67.4 95.5
LR 55.2 66.8 56.4 69.2 66.9 63.1 89.3
B 58.5 58.1 38.2 50.0 63.0 60.4 87.8
HSSD 28.5 34.5 47.2 0.0 47.8 20.0 50.0
Shortest path (Fréchet) K 56.4 55.8 53.5 82.6 66.0 66.1 94.4
LR 32.2 41.0 42.1 61.5 59.2 47.8 87.5
B 39.5 45.1 29.3 41.7 57.8 48.7 87.8
HSSD 22.5 27.3 13.9 0.0 34.8 20.7 50.0